### v2

# Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection

Abstract
Most toxicity detection models treat toxicity as an intrinsic property of text, overlooking the role of context in shaping its impact. Drawing on interdisciplinary research, we reconceptualise toxicity as a socially emergent stress signal. We introduce a new framework for toxicity detection, including a formal definition and metric, and validate our approach on a novel dataset, demonstrating improved contextual sensitivity and adaptability.
Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection
Sergey Berezin                        Reza Farahbakhsh SAMOVAR, Télécom SudParis, Institut Polytechnique de Paris, Palaiseau, France sergei.berezin@ip-paris.fr                        Noel Crespi
1 Introduction
Warning: This paper contains examples of language that may be perceived as offensive or toxic.
When asked to define obscenity, judge Potter Stewart famously said: “I know it when I see it". To this day, we have not come far from this definition.
Natural language processing (NLP) has made remarkable technical progress in detecting “toxic” language. Yet, despite ever-larger models and increasing benchmark scores, a foundational question remains unresolved: what exactly is toxicity?
An influential study by Dixon et al. (2018), defines toxic speech as “rude, disrespectful, or unreasonable language likely to make someone leave a discussion.” However, terms such as “rude” and “disrespectful” are highly subjective, varying considerably across cultures, individuals, and situational contexts. Crucially, such definitions offer no quantitative basis for measurement, leading to labelling inconsistencies that undermine model performance Garg et al. (2023). Despite all its vagueness and flaws, this definition is still widely used by researchers and industry.
Such ambiguity prevents consistent classification of even simple phrases - for instance, “I don’t like you”. Some may interpret it as toxic, others may not. Asking annotators with differing perspectives to label toxicity leads to disagreement, introducing noise that undermines model accuracy Liu et al. (2021); Goyal et al. (2022); Sap et al. (2021).
At the level of academia as a whole, the lack of a standardised definition has led researchers to employ different terms for similar work Vidgen (2019), assigning labels of “toxic", “hateful", “offensive", “abusive", etc. to effectively the same data Fortuna (2020); Madukwe et al. (2020). Considering this, in this paper, we use “toxic speech” as an umbrella term.
Another major issue in toxicity detection research is the field’s overreliance on benchmark-driven evaluation. Improvements in benchmark scores are frequently mistaken for real-world progress, despite the absence of ground truth metrics linking these scores to reductions in social harm such as exclusion, stress, or violence.
Most toxicity detection systems implicitly assume that toxicity is an intrinsic property of language - something that can be detected from surface features, isolated utterances, or predefined keyword lists. This assumption neglects the crucial role of context. Without context, toxicity detection systems are prone to racial, sexual, political, religious, and geographical biases Garg et al. (2023); Sap et al. (2019). These systems often degrade into simple profanity filters Chen (2022), lacking the nuance necessary for real-world use Xenos et al. (2021); Pavlopoulos et al. (2020); Menini et al. (2021) and becoming susceptible to numerous adversarial attacks Berezin et al. (2025, 2024, 2023).
Regarding the issues described above, researchers often argue that improving results requires a bigger and better model. Contrary to that, we suggest that we do not need a bigger model but a better definition.
1.1 Our Contribution: A New Frame, A Measurable Signal
We propose a paradigm shift: from viewing toxicity as an intrinsic property of language to understanding it as an emergent signal of perceived social stress. Drawing interdisciplinary research, we frame toxicity as a contextual stress reaction - one that becomes evident not solely through the content of a message, but through its social reception.
To operationalise this perspective, we introduce PONOS (Proportion Of Negative Observed Sentiments), a novel metric that quantifies toxicity based on the collective emotional response a message provokes. Rather than determining what a message intrinsically "is", PONOS captures how it is received.
2 Critics of current definitions
Contrary to the longstanding adage - "sticks and stones may break my bones, but words will never hurt me", contemporary research says: words can inflict measurable psychological and physiological harm Struiksma et al. (2022b); Sapolsky (2004).
Toxic communication has been linked to increased anxiety, chronic stress, radicalisation, and - at its extreme - self-harm and suicide van Geel et al. (2014); Kiritchenko et al. (2021); Cervone et al. (2021). At the societal level, hate speech exacerbates discrimination, fosters violence, and erodes the conditions for open discourse Vidgen and Derczynski (2021).
Yet despite the urgency of addressing these harms, the field lacks a coherent operational definition of toxicity. Definitions across academic, legal, and industrial contexts remain fragmented – often vague, inconsistent, and mutually incompatible Mnassri et al. (2024).
In an effort to systematise existing approaches, Herz and Molnár (2012) categorise definitions along four dimensions:
  1. 1.
    Harm-based: Speech that causes psychological or social harm;
  2. 2.
    Content-based: Speech conveying hateful ideas;
  3. 3.
    Lexical-based: Speech characterised by the use of derogatory words;
  4. 4.
    Dignity-based: Speech undermining the inherent dignity of target groups.
Each category, however, reveals critical shortcomings:
Harm-based definitions rely on an inherently vague notion of harm, which encompass both genuine toxicity and benign but distressing acts (e.g., delivering bad news).
Content-based definitions risk overreach, potentially classifying legitimate political disagreement as hate speech Parekh (2012); Hare and Weinstein (2009); Anderson and Barnes (2023).
Lexical-based approaches overlook toxic messages that use ostensibly neutral language (e.g., “Women, back to the kitchen!”).
Dignity-based models conflate legitimate critique with toxicity, risking the suppression of vital social and political discourse Anderson and Barnes (2023).
Subsequent work by Khurana et al. (2022) identifies five critical dimensions – target group, dominance, perpetrator identity, type of negative reference, and type of consequences – but finds no universal definition applicable across datasets or social contexts. Even expanded notions like Fiser et al.’s “socially unacceptable discourse” Fišer et al. (2017) fail to resolve the underlying contextual instability.
As Brown (2017) aptly observes, “hate speech” is not merely contested but “systematically ambiguous,” carrying different meanings depending on the cultural, political, and historical frames invoked.
3 Recontextualizing Toxicity: An Interdisciplinary Framework for Verbal Harm
To move beyond the limitations of current toxicity detection approaches, we synthesise insights from multiple fields to develop a framework for understanding toxic speech.
3.1 Personal level
The experience of verbal harm is fundamentally affective, rooted in how individuals perceive threats to personal and social stability. Research in psychology and neuroscience agrees that such harm does not stem from the intrinsic properties of language, but from context-dependent appraisals, emotional expectations, and violations of predictive frameworks.
In psychology appraisal theory Lazarus (1991) posits that emotional and stress responses arise when individuals evaluate events as threatening to their goals or social standing. In this view, communication becomes harmful not because of its content alone, but because it is perceived as violating significant personal concerns.
Research on social pain shows that symbolic threats can activate biological stress systems in much the same way as physical injury Eisenberger et al. (2003); Lieberman and Eisenberger (2009).
Neuroscientific findings indicate that the brain processes language in an anticipatory, predictive, and context-sensitive manner. Predictive coding models Friston (2010); Kleckner et al. (2017); Barrett and Simmons (2015) suggest that the brain constantly generates expectations about emotional and social states. When communication violates these expectations, it triggers stress responses.
Physiological indicators of stress, such as heart rate variability (HRV), are influenced by exposure to social threats Shaffer et al. (2014). ERP studies Ding et al. (2016); Liang et al. (2018) demonstrate that negative emotional language disrupts neural integration, while chronic exposure to social stressors leads to pro-inflammatory genetic activity Slavich and Cole (2013).
Together, this research shows that stress responses are closely tied to exposure to harmful or toxic speech, and that the resulting harm is not merely psychological - it is physiological as well.
3.2 Group level
Beyond direct victimization, witnessing harm to others can also trigger stress responses, signalling potential threats within the social environment Struiksma et al. (2022a).
Sociology and social psychology demonstrate that responses to verbal behaviour are governed by group norms, emotional expectations, and mechanisms of norm enforcement Elster (2007). Research by Bicchieri (2005) and Cialdini et al. (1991) shows that social norms are maintained through collective emotional and behavioural reactions to deviance. These violations are not assessed in isolation, but rather interpreted through the emotional dynamics of the group, which shape perceptions of harm. The concepts of emotional collectives van Kleef and Fischer (2015) and emotional norm-shaping processes van Kleef (2024); Sarı et al. (2024) emphasize how shared emotions reinforce or reshape social expectations in real time.
Moral Foundations Theory further underscores that emotional responses to norm violations differ across groups and cultures Graham et al. (2013)
From this perspective, toxicity functions as a social signal - a disruption in the moral and emotional fabric of the group.
3.3 Cultural-Language level
Linguistic and sociolinguistic research demonstrates that meaning is co-constructed in interaction, subject to contextual framing. Harmful speech cannot be reduced to individual words or expressions divorced from their social use.
Pragmatics Clark (1996); Sperber and Wilson (1986/1995) emphasises that meaning is not fully encoded in words but inferred through shared background knowledge and conversational context. An identical phrase may be benign in one setting and deeply offensive in another, depending on tone, power dynamics, and prior discourse.
Politeness theory and impoliteness theory Brown and Levinson (1987); Culpeper (1996) show that perceived rudeness, offence, or aggression are not intrinsic to utterances, but arise from how speech acts align (or misalign) with social expectations. In this view, offence is an emergent property of interaction — not a lexical attribute.
Language ideologies Woolard and Schieffelin (1994); Irvine (2001) highlight that judgements about what counts as “toxic” or “neutral” language are shaped by power structures and dominant cultural narratives.
Historical linguistics and semantics further demonstrates the change of meaning over the change of context Allan and Burridge (2006); Traugott and Dasher (2001). Words such as “nice” (originally meaning “foolish”) or “fag” (a benign term in British English, a slur in American usage) reveal how emotional and moral associations shift across cultural and historical contexts. This semantic fluidity undermines any fixed, universal standard of toxicity.
Multilingualism and language variation add further complexity. Emotional resonance and perceived offensiveness vary across languages due to differing socialisation patterns Dewaele (2004). What is taboo, rude, or inappropriate in one language and culture may carry no such connotation in another.
Together, these insights argue against treating harmful language as a static phenomenon. Toxicity cannot be reliably detected by identifying “bad words” in isolation, but only through an understanding of the forces that shape how language functions in context.
4 Definition of Toxicity
Building on those insights, we identify two core components of verbal harm: (1) context-dependent violation of social or moral norms, and (2) the induction of stress and emotional dysregulation as a consequence.
For example, consider a speaker who writes a profane, intentionally offensive message in Chinese to a recipient who does not speak the language. Although the content is objectively toxic and the intent is hostile, no harm is experienced if the recipient cannot interpret the message. There is no perceived norm violation, no affective response, and no stress. Under our framework, such a message would be classified as low in toxicity - not because it is morally defensible, but because it fails to generate a measurable signal of harm. The message would be toxic if it reached an informed audience.
Conversely, imagine a speaker makes a flirtatious joke, believing it to be benign or even complimentary. If the recipient experiences it as inappropriate or violating - due to context, power dynamics, or personal boundaries - stress and social disruption will occur. In many professional and legal contexts, this is recognised as harm: offence is taken, not given. Under our model, such an interaction would register as toxic, even in the absence of harmful intent.
These cases illustrate a crucial point: toxicity is defined not by intent or language alone, but by its reception - by the stress it induces and the norms it is perceived to violate.
Based on this synthesis, we offer the following definitions:
Definition 1.
Toxicity - the characteristic of causing stress through contradiction of accepted morality and norms of interaction within the context of interaction.
Consequently:
Definition 2.
Toxic speech - speech that causes stress through contradiction of accepted morality and norms of communication within the context of interaction.
These definitions deliberately foreground the role of context, perception, and reception. We distinguish between toxicity and toxic speech to account for the broader scope of harmful communicative acts, including visual, multimodal, or behavioural signals that may not be strictly linguistic.
It is important to clarify that while our framework uses stress as a measurable proxy for toxicity, we do not claim that all stress-inducing communication is inherently toxic. For instance, delivering bad news or asserting uncomfortable truths may elicit stress without violating social or moral norms. Our definition of toxicity focuses on the intersection of perceived norm violation and stress response, not on stress in isolation. This distinction is crucial to avoid conflating necessary discomfort with communicative harm.
This reconceptualisation challenges the assumption that toxicity can be located in isolated words or surface features. Instead, it positions toxicity as a socially emergent signal - a disruption to shared expectations that results in emotional or relational stress. Any detection system that hopes to mitigate real-world harm must take this interactional, context-sensitive perspective into account.
We are not denying intent’s relevance, but argue that it’s insufficient alone and serves rather as a part of a context. In our definition we use context as the full set of conditions under which a communicative act is interpreted - including verbal, situational, cognitive, cultural, and other socially relevant dimensions. The more precisely we understand these conditions, the more accurately we can assess the acceptability. A more detailed explanation can be found in Appendix A.
4.1 Towards Measurement: Stress as an Objective Metric
If the goal of toxicity detection is to mitigate harm, then we must define harm in terms that are both scientifically grounded and computationally tractable. Across disciplines, stress consistently emerges as a robust and measurable proxy for communicative harm, particularly in cases where harm is symbolic, ambiguous, or socially mediated.
Unlike subjective judgments of offensiveness or rudeness, stress is a quantifiable physiological state, akin to body temperature or blood pressure McEwen (1998); Thayer et al. (2012). It offers a principled anchor for modelling emotional impact, enabling us to move beyond vague or culturally biased notions of toxicity.
Stress can be inferred through a spectrum of physiological and behavioural indicators, each offering trade-offs between accuracy and scalability:
Direct physiological measures, such as glucocorticoid levels (e.g., cortisol), offer high biological accuracy and are widely used in laboratory studies. However, these require invasive sampling (blood, saliva, urine) and are not viable for real-time or large-scale deployment.
Non-invasive biosignals, including heart rate variability (HRV), galvanic skin response (GSR), pupil dilation, respiration, and facial or vocal affect, provide scalable approximations of stress. These signals correlate reliably with social and emotional threat responses Sharma and Gedeon (2012); Giannakakis et al. (2019), but still depend on hardware access and stable observation conditions.
Behavioural proxies, such as language use, reaction patterns, and discourse dynamics, offer the most feasible solution for large-scale deployment. In online settings, sentiment-laden replies, conversational shifts, and reaction memes often serve as collective indicators of stress. Prior research shows that these patterns reliably signal perceived norm violation and emotional disruption Jafari et al. (2023); Aleksandric et al. (2024).
5 PONOS: A Metric for Perceived Toxicity
Accepting toxicity being reflected in the evoked stress, we obtain observable proxy for stress levels. When a post elicits a wave of negative comments, disagreement, or distress, we see this as an indicator of perceived toxicity.
To formalise this, we introduce the Proportion Of Negative Observed Sentiments, abbreviated as PONOS (from Greek πόνος - "pain") — as an instantiation of our stress-based model. PONOS quantifies toxicity not through content analysis or classifier predictions, but through the emotional reception of a message within a specific community.
5.1 Defining PONOS
Let a post $x$ receive a set of replies $R(x)=\{r_{1},r_{2},\dots,r_{n}\}$. Each reply $r_{i}\in R(x)$ is assigned a sentiment score $s(r_{i})\in\{-1,0,+1\}$, where $-1$ denotes negative sentiment, 00 is neutral, and $+1$ is positive.
We define PONOS as the proportion of negative replies:
$\text{PONOS}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}\mathbb{I}[s(r_{i})=-1]$ (1)
where $\mathbb{I}[\cdot]$ is the indicator function.
This metric quantifies the perceived negativity of a post based on its community response. Higher values of $\text{PONOS}(x)$ indicate a stronger collective expression of disapproval or tension.
Compared to conventional approaches such as lexical sentiment counts, softmax-based toxicity classifiers, or black-box ML scores, PONOS metric offers the following advantages:
  • Interpretability: The value of PONOS directly reflects the proportion of negative replies, making it transparent and accessible to users, moderators, and downstream systems.
  • Context-relativeness: Because it reflects local community reactions, PONOS adapts to group norms, dialects, humour, and platform-specific conventions without requiring explicit encoding.
  • Statistical grounding: PONOS can be understood as the empirical mean of a Bernoulli distribution over negative/non-negative labels. By the law of large numbers, it converges to the true community-level disapproval rate as the number of replies increases — unlike softmax-based model scores, which often lack a consistent probabilistic interpretation (Guo et al., 2017).
  • Ratio-scale comparability: Because the score is ratio-scaled, posts with higher PONOS values are consistently more negatively received. For example, a score of 0.6 indicates twice the disapproval rate of 0.3. This enables meaningful ranking, thresholding, and aggregation — unlike many opaque ML toxicity scores, which cannot be meaningfully compared across models.
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context. In Appendix B we describe its variants, such as PONOS-Net, PONOS-Weighted, and PONOS-Early.
PONOS is not a universal measure of toxicity. It is a context-dependent metric — a community-grounded, reply-driven proxy for perceived social tension. Rather than determining what a message intrinsically "is", PONOS captures how it is received - reflecting the social impact of communication in its local context.
6 Empirical Evaluation
To demonstrate the viability of our framework for context-sensitive toxicity detection, we conducted a targeted empirical study. The process involves two main steps: first, identifying the most suitable sentiment function $s(\cdot)$ for the given context, and second, using it to select the best reaction prediction model for obtaining the PONOS value.
6.1 Dataset Creation
We curated a dataset from the Reddit community r/BlackPeopleTwitter. This community was chosen due to its large size (5.8 million members) and the fact that its guidelines explicitly encourage the use of African-American Vernacular English (AAVE) and slang.
Examples of AAVE often get misclassified as toxic speech by current toxicity detection models, which generally ignore contextual and cultural nuances. For example, a supportive phrase in AAVE like "That’s my nigga!" (roughly equivalent to “That’s my man!” in Standard American English) might be wrongly labelled as hateful or toxic. Such misclassifications effectively penalise speakers for communicating in their own dialect.
In total, we collected 401,931 posts containing 15,922,851 comments.
6.1.1 Annotation Process
To identify the optimal sentiment function $s(\cdot)$, we created a human-labeled test set from a subset of the data. We selected 56 user messages, each with at least four replies. If a message had too many replies, we retained only the top 7 by score, resulting in a total of 273 replies.
We conducted a structured labelling task using the Prolific platform. Annotators evaluated the sentiment of each reply, assigning two types of labels: (1) sentiment polarity — approval, neutral, or condemnation, and (2) sentiment target — whether the sentiment was directed at the original comment or the topic being discussed. This produced compound labels such as “approval (comment)” or “condemnation (subject).”
We recruited 11 annotators, all self-identifying as Black/African American U.S. nationals and regular internet users. The group included 7 men and 4 women, aged 21 to 60. To ensure high-quality data, we embedded attention checks and analysed labelling consistency. Four annotators showed strong engagement and reliable annotations; their labels were retained for further analysis.
We assessed inter-annotator agreement, obtaining a Fleiss’ Kappa of 0.48 and pairwise agreement of 67.8%, indicating moderately strong consensus given the subjectivity of the task. Final labels were determined using majority vote. In rare cases of a 2–2 tie, an additional annotation was provided by a researcher using the same labelling guidelines.
6.1.2 Ethical Considerations
To ensure our study adhered to ethical research standards, we took several measures. First, we complied with Reddit’s terms of service and ensured that no sensitive personal data was collected. We explained the nature of the task and data collection process to the participants beforehand, ensuring they were fully aware that the collected data would be used for research purposes.
Participants were compensated at a rate of \$12 per hour, in accordance with Prolific’s recommended payment guidelines. To estimate the time required for survey completion, we first completed the survey ourselves. The full text of the instructions given to participants can be found in Appendix D.
6.2 Experiments
We conducted all experiments using NVIDIA H100 GPUs, totaling 883 compute hours. Model training and evaluation were performed using the February 2025 release of the Unsloth library (unslothai, 2025). Metrics were computed using scikit-learn version 1.5.0 Pedregosa et al. (2011).
We were committed to providing as much context as possible to the models, and we found this to be beneficial for the models’ performance. The input data for sentiment classification of a reply to a comment included: the comment itself, the text of all replies for classification, the title of the post. The prompt used with LLMs is provided in Appendix E.
6.2.1 Sentiment Analysis Model
To determine the most suitable sentiment analysis function $s(\cdot)$ for use in Equation 1, we evaluated both specialized sentiment classifiers and general-purpose LLMs using custom prompts. Evaluation was conducted on the human-labelled subset of our dataset. Results are presented in Table 1.
Model Prec. Rec. F1
RoBERTa (2023) 0.63 0.22 0.32
ModernBERT (2024) 0.71 0.21 0.32
Twitter-RoBERTa (2020) 0.58 0.64 0.61
GPT-4o (2024a) 0.67 0.82 0.74
LLaMA 3 70B 0.81 0.70 0.75
Qwen 3 70B 0.76 0.80 0.78
Table 1: Precision, recall, and F1 scores for negative sentiment class across various models.
Qwen 3 Yang et al. (2025) showed the best overall performance and was selected for all subsequent experiments. During error analysis, we found that most misclassifications occurred in examples with low inter-annotator agreement. These instances often involved ambiguous language that led both assessors and models to varied interpretations.
6.2.2 Toxicity Prediction
With a sentiment analysis model selected, we computed PONOS scores for each post in the dataset. To evaluate binary toxicity classification, we thresholded the continuous PONOS values: posts with $\text{PONOS}\geq 0.5$ were labeled as toxic, and all others as non-toxic. The threshold of 0.5 reflects the intuition that at least half of readers are likely to respond negatively. It also aligns with thresholds used in many existing moderation frameworks, facilitating fair comparison with baseline models.
We made the following evaluations:
Direct PONOS Prediction:
A regression model was trained to predict the gold-standard PONOS score from post text alone. This method is simple and computationally efficient, making it suitable for real-time toxicity detection using lightweight models.
PONOS Prediction via Simulated Reactions:
A language model was trained to generate plausible replies to a given post. These synthetic replies were passed through the sentiment analysis model to estimate the PONOS score. This method aligns with our theoretical framing of toxicity as an emergent social signal: it approximates how a community might react.
Comparison with Baselines:
As baselines, we include predictions from the OpenAI Moderation API (omni-moderation-latest) (2024b), Google Perspective API (2024), and Detoxify (2020). These systems represent current industry standards and enable comparison in terms of contextual sensitivity and error types.
Ablation Study:
To isolate the contribution of domain-specific adaptation and reaction-modeling strategies, we conducted an ablation study using LLaMA 3 70B Llama Team (2024) across four variants:
  • TP (Toxicity Prediction): Predicts binary toxicity label directly from the input utterance.
  • DTTP (Domain-Tuned Toxicity Prediction): Same as TP, but with domain-specific pretraining on the unlabeled portion of our dataset.
  • RP (Reply Prediction): Generates replies to the input. Toxicity is then inferred by applying a thresholded PONOS score computed from the predicted replies using a sentiment analysis model.
  • DTRP (Domain-Tuned Reply Prediction): Same as RP, but with domain-specific pretraining on the unlabeled dataset.
6.2.3 Large Language Model Pretraining
In DTTP and DTRP variants we pretrained a LLaMA 3 70B model using domain-specific Reddit data from r/BlackPeopleTwitter. The objective was to perform domain adaptation, enabling the model to generate plausible, community-authentic replies (DTRP) or contextually-aware toxicity prediction (DTTP).
Model Architecture and Configuration
We used the unsloth framework to load and fine-tune the base model in 4-bit precision. The model was adapted using LoRA with rank 64 for efficient parameter updates. Gradient checkpointing and rslora were employed to further reduce memory overhead. All training and inference were conducted using bfloat16 precision and FlashAttention 2 to maximize memory efficiency.
Prompt Format and Dataset Construction
Each training example consisted of a Reddit comment followed by a prompt requesting five likely replies in the tone and style of r/BlackPeopleTwitter. Replies were separated by a special delimiter token <|reply|>. The prompt template is provided in Appendix E.
Model MSE MAE
Direct Prediction
TF-IDF + Ridge 0.046 0.173
SentenceBERT FT 0.043 0.166
Reply Prediction
LLaMA 3 8B 0.278 0.402
LLaMA 3 70B RP 0.138 0.304
GPT-4o 0.214 0.308
LLaMA 3 70B DTRP 0.042 0.152
Table 2: Mean Squared Error (MSE) and Mean Absolute Error (MAE) for PONOS score prediction . FT = finetuned model.
Model Prec. Rec. F1
Direct Prediction
TF-IDF + Ridge 0.63 0.26 0.36
all-mpnet + Ridge 0.65 0.30 0.41
SentenceBERT FT 0.54 0.65 0.59
Pretrained Baselines
OpenAI moderation 0.43 0.28 0.34
Perspective API 0.35 0.31 0.33
Detoxify 0.35 0.31 0.33
Detoxify-unbiased 0.35 0.34 0.35
Reply Prediction and Ablation Study
LLaMA 3 70B TP 0.43 0.40 0.42
LLaMA 3 70B DTTP 0.42 0.42 0.42
LLaMA 3 70B RP 0.40 0.21 0.27
LLaMA 3 70B DTRP 0.67 0.71 0.69
Table 3: Precision, recall, and F1 score for toxicity classification (toxic class only) across direct prediction, pretrained baseline, and reply-based models.
6.3 Results
As shown in Tables 3 and 2, our best model - DTRP (Domain-Tuned Reply Prediction), achieves a substantial improvement in F1 score (0.69), significantly outperforming both direct PONOS predictors (e.g., finetuned SentenceBERT (2019) at 0.59) and external moderation tools such as Detoxify, OpenAI Moderation API, and Perspective API (F1 $\approx$ 0.33–0.35). This confirms the core hypothesis of our work: toxicity is better modelled as contextual social friction than as a fixed textual property.
Ablation Study further illustrates that it’s the combination of domain adaptation and reply-based modelling show the greatest gains.
On top of that, simulated reactions offer several advantages: they incorporate social and cultural context, allow integration of multimodal input (via multimodal LLMs), and provide interpretability by exposing intermediate responses. This approach also generalizes well to novel or underspecified content.
Complementing these quantitative results, our error analysis (Appendix C) reveals the systemic limitations of existing models.
Commercial tools frequently miss ideologically disruptive or passive-aggressive content (false negatives), while overflagging vernacular, humor, and identity expression (false positives). These failures reflect an overreliance on surface features and global thresholds. By contrast, PONOS captures perceived harm through community reactions, allowing our models to align more closely with socially grounded definitions of toxicity.
7 Conclusion: Toward a Contextual Ethics of Harm
Toxicity is not an intrinsic property of language, but an emergent social signal - a reaction to perceived norm violation and induced stress within a particular context. In this framing, harm is not universal but situational; not static but dynamic. What matters is not what is said in isolation, but how it is received and interpreted.
By reframing toxicity detection as the measurement of perceived social friction, we challenge NLP’s prevailing paradigm of lexical essentialism. Instead, we propose a framework grounded in psychology, neuroscience, and computational social science.
Our metric, PONOS, embodies this shift. Rather than classifying messages based on intrinsic properties, PONOS captures their collective social reception. It is interpretable, context-sensitive, and empirically grounded - a pragmatic tool for modelling harm as it is experienced by real communities.
PONOS is not a moral arbiter. It does not tell us what is right, but what is stressful. It is a mirror, not a judge. Moderation systems should treat PONOS as an advisory signal, not a censorship trigger. High PONOS scores do not necessarily indicate ethical failure; they signal social dissonance. Low scores are not always desirable either - total agreement may reflect echo chambers or silencing.
We are not offering a final definition of toxicity, but a new conceptual framework for studying it - one that acknowledges context, perception, and affect. PONOS illustrates that subjectivity can be modelled without being reduced.
Future work can extend this approach by improving the model, incorporating multimodal signals, conducting cross-cultural comparisons, and validating it in real-world scenarios.
Ultimately, we aim to shift NLP from classifying bad words to understanding social dynamics - from linguistic essentialism to social contextualism.
8 Limitations
We acknowledge several important limitations of our framework:
  • Stress is not synonymous with toxicity. Not all norm violations are harmful (e.g., satire, protest), and not all harm produces measurable stress. Our metric captures perceived disruption, not ethical correctness. It defines toxicity at the intersection of norm violation and stress response — not stress alone — to avoid conflating necessary discomfort with communicative harm.
  • Sentiment classification is an imperfect proxy. PONOS relies on sentiment classifiers that are known to carry cultural, demographic, and linguistic biases. Their subjectivity introduces noise, especially in edge cases like sarcasm or coded language.
  • Community-level perception $\neq$ moral truth. PONOS reflects what provokes disapproval in a particular group, not a universal judgment. What is toxic in one community may be normative in another. This is a strength - but also a limitation.
  • Susceptibility to manipulation. Like all crowd-based signals, PONOS can be gamed through brigading, reply injection, or sockpuppetry. Mitigating these risks may require extensions such as reply weighting (e.g., verified users, community trust scores, vulnerability-aware weighting) or time-window filters.
  • Blind spots. PONOS cannot capture silent harm - for example, when content is ignored, downvoted, or drives users away without leaving replies. It may miss the damage caused by exclusion or indirect effects.
  • Not a moral compass. PONOS is not a substitute for ethical judgment. It models perception, not morality. It should be used alongside normative frameworks, not in place of them.
Despite these caveats, we argue that PONOS offers the most scalable, interpretable, and behaviourally grounded operational proxy for communicative harm currently available - especially for real-world NLP systems.
9 Ethical Considerations
Any metric used in content moderation, filtering, or prioritization carries normative weight. While we position PONOS as a descriptive tool - reflecting community reactions rather than imposing moral standards - we recognize its potential for misuse or overreach.
Moderation decisions based solely on PONOS scores risk reinforcing existing biases or silencing dissent. High PONOS should not automatically lead to suppression; rather, it should signal the need for human oversight, context analysis, and perhaps dialogic intervention.
We explicitly do not claim that PONOS reflects objective truth. It reflects perceived social friction - a volatile signal that can be shaped by group norms, visibility effects, and adversarial influence.
Instead, we call for responsible deployment:
  • Use PONOS transparently and explainably.
  • Pair it with community input, human review, and ethical guidance.
  • Monitor for drift, manipulation, and misalignment with social values.
  • Consider its use not just for suppression, but also for norm diagnosis, echo chamber detection, and semantic change tracking.
Ultimately, we advocate for a shift in perspective: from minimizing toxicity to mapping social tension. This enables more pluralistic, adaptive, and humane approaches to language technology - aligned not with abstract ideals, but with the lived experience of language in context.
Appendix A Details on Understanding of Context
Context structures interpretation. It governs what is considered appropriate, offensive, or harmful within a given setting. To support operational clarity, we distinguish four key dimensions:
Verbal context refers to the surrounding linguistic material — what is said before, after, or alongside the utterance. It determines how meaning is disambiguated, particularly in cases of sarcasm, reappropriation, or coded language.
Situational context includes the physical, social, and institutional setting of communication: who is speaking, to whom, in what place, and under what conditions. Normative expectations can vary dramatically between settings — for example, language deemed acceptable at a music festival may be inappropriate in a courtroom or classroom. This dimension also encompasses digital and symbolic environments such as online platforms, workplace hierarchies, or subcultural spaces.
Cognitive context refers to the mental and emotional states of the communicative participants — including intentions, assumptions, prior knowledge, and emotional stance. Many speech acts, such as jokes, flirtation, or irony, can only be correctly interpreted when cognitive context is inferred or shared.
Cultural context involves the broader moral frameworks, communicative conventions, and ideological norms that vary across communities. A behaviour considered respectful in one culture may be viewed as inappropriate or even threatening in another. For example, direct individualistic expression is normative in many Western contexts, but may be perceived as disruptive or disrespectful in collectivist settings.
Given this multidimensional variability, any model of toxicity detection must be contextually grounded and culturally adaptive. Static or universalist rule sets will inevitably misclassify utterances as they fail to consider how meaning emerges within interaction.
Appendix B PONOS Variants
B.1 Extensions and Variants
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context.
PONOS-Net.
We define the net sentiment variant as:
$\text{PONOS}_{\text{net}}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}s(r_{i})$ (2)
This version captures the overall affective balance of replies, distinguishing polarized threads (with both positive and negative responses) from uniformly negative ones.
PONOS-Weighted.
To account for user-level features, such as author credibility, social influence, or vulnerability - we define a weighted version:
$\begin{split}\text{PONOS}_{\text{w}}(x)=\frac{1}{\sum_{i}w(r_{i})}\sum_{r_{i}\in R(x)}w(r_{i})\cdot\mathbb{I}[s(r_{i})=-1]\end{split}$ (3)
Here, $w(r_{i})$ is a weight assigned to each reply, reflecting community-specific salience (e.g., upvotes, karma, user type).
PONOS-Early.
To mitigate feedback loop effects (e.g., reply priming), we define an early-window variant that considers only the first $k\ll n$ replies:
$\text{PONOS}_{\text{early}}(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}[s(r_{i})=-1]$ (4)
This variant reflects the immediate, unprimed perception of a post, prior to potential community escalation or reinforcement.
B.2 Implementation Considerations
The sentiment function $s(\cdot)$ can be realized through multiple approaches:
  • Manual annotation by culturally aligned raters;
  • Off-the-shelf sentiment classifiers (e.g., RoBERTa, BERTweet, GPT-based);
  • Community-specific fine-tuned models.
We emphasize that the interpretability and validity of PONOS depend on the alignment between the sentiment model and the communicative norms of the target community. In our experiments, we rely on human-annotated sentiment labels sourced from subreddit-native annotators to ensure cultural fidelity.
Appendix C Error Analysis of Moderation Tools
To diagnose systemic weaknesses in current toxicity classifiers, we conducted an error analysis focused on two extreme categories: absolute false negatives (PONOS = 1.0, but flagged as non-toxic by all baselines) and absolute false positives (PONOS = 0.0, but flagged as toxic by all baselines). These cases reveal the most severe misalignments between models’ predictions and actual community perception.
False Negatives: Missed Norm Violations
Posts with PONOS = 1.0 consistently provoked hostile, aggressive, or sarcastic replies, signalling strong perceived norm violations. Yet commercial tools such as OpenAI’s moderation API and Google’s Perspective API assigned very low toxicity probabilities (mean scores of 0.089 and 0.159, respectively), failing to detect any issue.
Common patterns include:
  • Sarcastic antagonism — Politically or ideologically provocative statements framed politely. For instance, "Employees shouldn’t be paid proportionally to profit." triggered strong public disapproval, despite lacking overtly toxic language.
  • Tone-dependent offences — Subtle or veiled hostility that eludes profanity filters and shallow classifiers.
Only 4% of these false negatives contained any profanity, highlighting that keyword-reliant models systematically overlook covert or contextually-coded toxicity.
False Positives: Overflagging Harmless Content
Conversely, posts with PONOS = 0.0 - indicating no perceived harm by the community - were frequently over-flagged by moderation models (mean OpenAI score = 0.76; Perspective API = 0.70).
Failure patterns include:
  • Misclassified AAVE - Community-accepted slang such as "hood niggas are some of the biggest nerds…"; Y’all niggas remember Neopets tho?" was flagged despite being clearly humorous and affectionate.
  • Penalized identity celebration - Comments expressing racial pride or solidarity (e.g., "Love being Black, but I gotta stop you on the Black church…") were frequently overflagged.
  • Jokes and playful tone - Humorous remarks like "I, too, am happier when her titties are out." (referencing a celebrity quote about not wearing a bra) were misinterpreted as harmful.
While profanity appeared in 62% of false positives, contextual cues and response sentiment clearly indicated benign usage, revealing the brittleness of static filters.
Thematic Breakdown
Topic modelling showed that both error types disproportionately involved posts about race, gender, and politics. However, the failures diverged in direction: False negatives frequently involved ideological friction with minimal keyword salience; False positives were dominated by race-related humour or cultural expression.
Appendix D Assessors’ Instructions
Dear Participant,
Thank you for taking the time to participate in our survey. Your input is invaluable to us. In this survey, you will be asked to classify reactions expressed in replies to comments from discussions on the Reddit community.
Task:
- Given a comment and its replies from a discussion on the Reddit community r/BlackPeopleTwitter, classify the reaction expressed in each reply as one of the following:
- "approval (comment)" - Agreement/support of the comment itself.
- "approval (subject)" - Agreement/support of the subject mentioned in the comment.
- "neutral" - No strong reaction for or against the comment or subject.
- "condemnation (comment)" - Disapproval of the comment itself.
- "condemnation (subject)" - Disapproval of the subject mentioned in the comment.
- Be sure to identify whether approval/condemnation is directed at the comment or the subject.
Examples:
Topic: "New office policy - no phones allowed"
Comment: "The new policy is a disaster, its making everything worse."
1. Reply: "I totally agree, its a complete mess."
- Classification: approval (comment)
2. Reply: "I think we should give it some time to see the effects."
- Classification: neutral
3. Reply: "This policy is the worst decision ever made by this administration."
- Classification: condemnation (subject)
4. Reply: "I dont have a strong opinion on this."
- Classification: neutral
5. Reply: "LOL, shut up Karen, nobody cares what you think."
- Classification: condemnation (comment)
6. Reply: "I kinda like the new policy."
- Classification: approval (subject)
By following these instructions, you will help us understand the community dynamics and how different reactions are expressed within the r/BlackPeopleTwitter community.
Your participation is crucial for the success of this research. Thank you!
Appendix E Prompts
Sentiment classification prompt:
You are a sentiment analysis assistant trained to understand African American Vernacular English (AAVE) and online Black community discourse.
Your task is to classify the emotional tone of REPLY message as Positive, Neutral, or Negative.
Your analysis supports a larger framework for detecting *perceived toxicity* - understanding how it might cause stress or social tension within the community.
Use all available context:
- Topic: the post that started the thread
- Comment: the message being replied to
- Reply: the specific message to analyze
Consider slang, tone, cultural style, and conversational intent.
- Positive: Agreement, Pride, joy, affirmation, playfulness, flirting, or supportive sarcasm. Being angry TOGETHER against something also goes here.
- Neutral: Factual comments, observations, dry or ambiguous tone, rhetorical questions.
- Negative: Disagreement, confrontation, arguing, complaints, passive aggression, warnings, critique, or social "shade."
Always output one of: Positive’, Neutral’, or Negative without explanation. Be sure to not miss any Negative’.
Format your answer as "ANSWER:"
FEW-SHOT EXAMPLES:
Topic: Trump won
Comment:" I think this is ok as long as we are not getting democrat president
Reply: Nigga this is a HILARIOUS take
ANSWER: Negative
Topic: New record lol
Comment: That clip had me dying, fr.
Reply: Ayo I been laughing for 10 minutes, why nobody brings up that he claimed to be THE MASTER
ANSWER: Positive
Topic: Gotta do whacha gotta do
Comment: Once I was texting with girl for 6 months to get a picture from her
Reply: Bro, Ive ended up writing her real letters. desparate shit
ANSWER: Neutral
Topic: Economy is fucked
Comment: I finna strat looting just to get food on the table
Reply: they are making us do it
ANSWER: Positive
Reaction prediction prompt:
{"role": "system", "content": "You are a culturally aware assistant who creates authentic replies from r/BlackPeopleTwitter."},
{"role": "user", "content": (
"Below is a Reddit comment. Generate the top 5 likely replies from r/BlackPeopleTwitter, "
[…]