### v3

# From Intrinsic Toxicity to Reception-Based Toxicity: A Contextual Framework for Prediction and Evaluation

Abstract
Most toxicity detection models treat toxicity as an intrinsic property of text, overlooking the role of context in shaping its impact. In this position paper, drawing on insights from psychology, neuroscience, and computational social science, we reconceptualise toxicity as a socially emergent signal of stress. We formalise this perspective in the Contextual Stress Framework (CSF), which defines toxicity as a stress-inducing norm violation within a given context and introduces an additional dimension for toxicity detection. As one possible realisation of CSF, we introduce PONOS (Proportion Of Negative Observed Sentiments), a metric that quantifies toxicity through collective social reception rather than lexical features. We validate this approach on a novel dataset, demonstrating improved contextual sensitivity and adaptability when used alongside existing models.
From Intrinsic Toxicity to Reception-Based Toxicity: A Contextual Framework for Prediction and Evaluation
Sergey Berezin          Reza Farahbakhsh SAMOVAR, Télécom SudParis, Institut Polytechnique de Paris, Palaiseau, France sergei.berezin@ip-paris.fr          Noel Crespi
1 Introduction
Warning: This paper contains examples of language that may be perceived as offensive or toxic.
When asked to define obscenity, Judge Potter Stewart famously said: "I know it when I see it". To this day, we have not come far beyond this intuition.
Natural language processing (NLP) has made remarkable technical progress in detecting "toxic" language. Yet, despite ever-larger models and rising benchmark scores, a foundational question remains unresolved: what exactly is toxicity?
An influential work by def_old defines toxic speech as "rude, disrespectful, or unreasonable language likely to make someone leave a discussion". However, terms such as "rude" and "disrespectful" are inherently subjective, varying considerably across cultures, individuals, and situational contexts. Crucially, such definitions offer no quantitative basis for measurement, leading to labelling inconsistencies that undermine model performance survey.
Despite its vagueness and limitations, this definition remains widely used in both research and industry.
This ambiguity prevents consistent classification even for seemingly simple phrases - for instance, "I don’t like you". Some may interpret it as toxic, while others may not. Asking annotators with differing perspectives to label toxicity therefore leads to disagreement, introducing noise that undermines model accuracy authors1; autors2; authors3.
At the level of academia as a whole, the lack of a standardised definition has led researchers to employ different terms for similar phenomena vidgen-etal-2019-challenges, assigning labels such as "toxic", "hateful", "offensive", or "abusive" to effectively the same data its_all_same; madukwe-etal-2020-data. In this paper, we use toxic speech as an umbrella term.
Another major issue in toxicity detection research is the field’s overreliance on benchmark-driven evaluation. Improvements in benchmark scores are frequently treated as evidence of real-world progress, despite the absence of ground-truth metrics linking these scores to reductions in social harm, such as exclusion, stress, or violence.
Most toxicity detection systems implicitly assume that toxicity is an intrinsic property of language - something that can be detected from surface features, isolated utterances, or predefined keyword lists. This assumption neglects the crucial role of context. Without context, toxicity detection systems are prone to racial, sexual, political, religious, and geographical biases survey; bias_nig. These systems often degrade into simple profanity filters holy_shit, lacking the nuance necessary for real-world use xenos-etal-2021-context; context2; context3 and becoming susceptible to numerous adversarial attacks tip; berezin2024; berezin-etal-2023-offence.
1.1 Our Contribution: A New Frame, A Measurable Signal
To address this, we argue for a shift in emphasis: from viewing toxicity as purely intrinsic to the text towards modelling it as an emergent signal of perceived social stress.
Drawing on interdisciplinary insights, we introduce the Contextual Stress Framework (CSF), which frames toxicity as a stress-inducing norm violation within a given context. CSF adds an additional dimension to existing text-intrinsic approaches by capturing how toxicity is perceived and socially received.
Within CSF, we introduce PONOS (Proportion Of Negative Observed Sentiments) as one possible realisation of the framework: a metric that quantifies toxicity via the collective emotional reception a message provokes. Rather than determining what a message intrinsically is, PONOS captures how it is received in context.
2 Related Work and Limitations of Existing Definitions
Contemporary research demonstrates that words can inflict measurable psychological and physiological harm struiksma2022people; sapolsky2004zebras.
Toxic communication has been linked to increased anxiety, chronic stress, radicalisation, and, at its extreme, self-harm and suicide VanGeel2014; Kiritchenko2021; Cervone2021. At the societal level, hate speech exacerbates discrimination, fosters violence, and erodes the conditions for open public discourse Vidgen2020.
Despite the urgency of addressing these harms, the field lacks a coherent operational definition of toxicity. Definitions across academic, legal, and industrial contexts remain fragmented, often vague, inconsistent, and mutually incompatible mnassri2024survey.
In an effort to systematise existing approaches, Herz2012 categorise definitions of harmful speech along four dimensions:
  1. 1.
    Harm-based: speech that causes psychological or social harm;
  2. 2.
    Content-based: speech conveying hateful ideas;
  3. 3.
    Lexical-based: speech characterised by the use of derogatory words;
  4. 4.
    Dignity-based: speech undermining the inherent dignity of target groups.
Each category, however, exhibits critical limitations.
Harm-based definitions rely on an inherently underspecified notion of harm, encompassing both genuinely toxic speech and socially necessary but distressing acts (e.g. delivering bad news).
Content-based definitions risk overreach, potentially classifying legitimate political disagreement or critique as hate speech parekh_2012; hare; sep-hate-speech.
Lexical-based approaches overlook toxic messages expressed through ostensibly neutral language (e.g. "Women, back to the kitchen!").
Dignity-based models risk conflating legitimate criticism with toxicity, thereby suppressing vital social and political discourse sep-hate-speech.
Subsequent work by khurana-etal-2022-hate identifies five critical dimensions: target group, dominance, perpetrator identity, type of negative reference, and type of consequences. Yet, it finds no universal definition applicable across datasets or social contexts. Even broader formulations, such as Fiser et al.’s notion of "socially unacceptable discourse" fiser, fail to resolve the underlying contextual instability.
As Brown2017 observes, harmful speech is not merely contested but systematically ambiguous, acquiring different meanings depending on the cultural, political, and historical frames in which it is situated.
3 An Interdisciplinary Basis for CSF
Across psychology, neuroscience, and sociolinguistics, verbal harm is understood as fundamentally context-dependent: the same utterance can be harmless or harmful depending on who interprets it, under what expectations, and within which normative setting. A common thread is an appraisal process: communicative acts are evaluated relative to local moral and interactional norms, and when an act is perceived as threatening, disrespectful, or boundary-violating, it can trigger affective responses and stress-related dysregulation in recipients and bystanders lazarus1991emotion; eisenberger2003does; lieberman2009neuroscience.
At the individual level, research on stress and emotion emphasises that harmful impact depends on interpretation rather than surface form alone. Predictive and context-sensitive accounts of cognition model perception as expectation-driven; violations of anticipated social or emotional states can elicit stress responses even when intent is ambiguous friston2010free; barrett2015interoceptive; kleckner2017evidence. This link between social threat and physiological regulation is reflected in established correlates such as heart-rate variability shaffer2014healthy.
At the interactional level, harm is jointly constructed and becomes legible through sequences of action and reaction: norms are maintained and renegotiated through observable responses such as condemnation, withdrawal, or escalation. Work on norm enforcement shows that deviance is regulated through collective emotional and behavioural reactions Bicchieri_2005; CIALDINI1991201; vankleef2015emotions.
At the cultural level, communities differ in how they sanction profanity, sarcasm, reclaimed slurs, or in-group teasing; what appears "toxic" out of context may be benign in-group signalling, and vice versa. Pragmatic accounts emphasise that meaning is inferred from shared background and interactional framing rather than encoded in words alone clark1996using; Sperber1995-SPER. Politeness and impoliteness research likewise treat offence as an emergent property of interaction Brown_Levinson_1987; CULPEPER1996349, while language ideology work highlights how judgements of "toxic" versus "neutral" language are shaped by power and dominant cultural narratives woolard1994language; irvine2001style. Finally, meaning and offensiveness shift across time and across languages, undermining any fixed universal standard Traugott_Dasher_2001; Dewaele01062004.
Taken together, these literatures motivate CSF’s core composition: toxicity arises when an utterance is perceived to violate local norms and this perception induces stress or disruption in an audience (additional discussion in Appendix A).
4 Definition of Toxicity
Building on the interdisciplinary insights outlined above, we model verbal harm as comprising two components: (i) a perceived violation of social or moral norms in a given interactional context; and (ii) an ensuing stress response in those who interpret the act.
Consider a speaker who writes a profane, intentionally offensive message in Chinese to a recipient who does not speak the language. Although the intent is hostile and the string contains profanity, no harm is experienced if the recipient cannot interpret it. There is no perceived norm violation, no affective response, and no stress. Under our framework, such a message registers as low in toxicity in this interaction - not because it is morally defensible, but because it fails to produce a measurable signal of harm in the audience that actually receives it. The same message would become toxic if it reached an informed audience.
Conversely, imagine a speaker makes a flirtatious joke believing it to be benign or even complimentary. If the recipient experiences it as inappropriate or violating - due to context, power dynamics, or personal boundaries, stress and social disruption may occur. In many professional and legal contexts, this is recognised as harm: offence is taken, not given. Under our model, such an interaction can be toxic even in the absence of harmful intent.
These examples motivate a crucial point: toxicity is defined not by intent or language alone, but by its reception - by the stress it induces and the norms it is perceived to violate.
Based on this synthesis, we propose:
Definition 1.
Toxicity — the characteristic of causing stress via a perceived contradiction of accepted morality and norms of interaction, within the context in which it is interpreted.
Consequently:
Definition 2.
Toxic speech — speech that induces stress via a perceived contradiction of accepted morality and norms of communication, within the context in which it is interpreted.
These definitions deliberately foreground context, perception, and reception. We distinguish between toxicity and toxic speech to account for the broader scope of harmful communicative acts, including visual, multimodal, or behavioural signals that may not be strictly linguistic.
Operational implication (two-axis view). The two components of the definition above yield two complementary measurement routes. Text-only toxicity detectors (e.g. moderation models trained on annotator judgements) primarily operationalise perceived norm violation from the content itself under assumed or averaged contexts. In contrast, reception-based signals operationalise experienced stress as it manifests in interaction (e.g. in replies), capturing context that is not present in the message text alone. Our framework therefore treats intrinsic scores and reception signals as distinct axes rather than interchangeable "toxicity scores": an utterance may be norm-violating without eliciting strong negative reception (high violation, low stress), or elicit negative reception without being flagged by text-only detectors (low violation, high stress).
While our framework treats stress as a measurable proxy for harm, we do not claim that all stress-inducing communication is inherently toxic. Delivering bad news or asserting uncomfortable truths may elicit stress without violating social or moral norms. Toxicity, as defined here, arises from the intersection of perceived norm violation and stress response, not from stress in isolation. This distinction is crucial to avoid conflating necessary discomfort with communicative harm.
This reconceptualisation challenges the assumption that toxicity can be located in isolated words or surface features. Instead, it positions toxicity as a socially emergent signal - a disruption of shared expectations that manifests as emotional or relational stress. Any detection system intended to mitigate real-world harm must therefore adopt an interactional, context-sensitive perspective.
We do not deny the relevance of speaker intent, but argue that intent alone is insufficient and should instead be treated as one component of context. Here, context refers to the full set of conditions under which a communicative act is interpreted, including linguistic, situational, cognitive, cultural, and relational dimensions. A more detailed discussion of contextual factors is provided in Appendix B.
4.1 Towards Measurement: Stress as a Measurable Signal
If toxicity detection aims to mitigate harm, then the relevant notion of harm must be grounded in a signal that is (i) theoretically motivated and (ii) measurable at scale. Across disciplines, stress provides such an anchor: it refers to a family of physiological and behavioural responses that arise when an individual appraises a situation as threatening, boundary-violating, or socially unsafe. This makes stress a principled correlate of communicative harm in cases where harm is symbolic, ambiguous, or socially mediated.
Unlike global labels such as "offensive" or "rude", stress has a well-characterised biological basis and a long measurement tradition in psychology and neuroscience mcewen1998; thayer2012.
Stress can be estimated through multiple classes of indicators, each trading off fidelity against scalability. Direct physiological measures (e.g. cortisol levels) are biologically faithful but require invasive sampling and are impractical outside controlled settings. Non-invasive biosignals such as heart-rate variability, galvanic skin response, respiration, pupil dilation, and facial or vocal affect provide scalable approximations and correlate with social-threat responses sharma2012objective; soc_stress_biosignals, but still depend on specialised hardware and stable observation conditions.
For NLP systems deployed in online environments, the most practical evidence is therefore behavioural: stress and disruption often surface in interaction through reaction patterns, shifts in tone, escalation, withdrawal, or affective language in replies. Prior work shows that collective response dynamics can serve as indirect indicators of perceived norm violation and emotional disruption in online discussion Amir; Aleksandric_2024. Importantly, behavioural proxies do not measure internal physiology directly; they capture how social tension manifests publicly in interaction - precisely the level at which toxicity becomes consequential for communities and moderation decisions.
5 PONOS: A Metric for Reception-Based Toxicity
Building on the stress-based account of toxicity developed above, we now introduce a concrete metric for capturing reception-based toxicity: how a communicative act is received by its audience in context.
Our definition implies a causal pathway: a post is toxic when it is perceived to contradict local moral or interactional norms, and this perceived violation induces a stress response in the audience.
In online interaction, stress and social disruption often surface as expressions of disapproval, outrage, distress, or condemnation in replies. Accepting that toxicity is reflected in the stress it evokes, we treat patterns of community response as a proxy signal for reception-based toxicity: when a post elicits a disproportionate share of negative reactions, this provides evidence that it violated local norms and induced tension in its audience.
To formalise this intuition, we introduce PONOS (Proportion Of Negative Observed Sentiment; from Greek πόνος, "pain") as one instantiation of the Contextual Stress Framework. PONOS quantifies toxicity not through lexical content or opaque model scores, but through the emotional reception of a message within a specific community and context.
5.1 Defining PONOS
Let a post $x$ receive a set of replies $R(x)=\{r_{1},r_{2},\dots,r_{n}\}$. Each reply $r_{i}\in R(x)$ is assigned a sentiment label $s(r_{i})\in\{-1,0,+1\}$, where 1-1 denotes negative sentiment, 0 neutral sentiment, and +1+1 positive sentiment.
We define PONOS as the proportion of negative replies:
$\text{PONOS}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}\mathbb{I}[s(r_{i})=-1],$ (1)
where $\mathbb{I}[\cdot]$ denotes the indicator function.
Intuitively, $\text{PONOS}(x)$ estimates the share of observed replies that express negative affect. Higher values indicate stronger collective expressions of disapproval or distress, consistent with reception-based harm in context.
PONOS is (i) interpretable as a proportion of negative replies, (ii) context-sensitive via local audience norms, and (iii) statistically grounded and comparable, as an estimate of a negative-reception rate whose uncertainty decreases with $|R(x)|$ (guo2017calibration).
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context. In Appendix C we describe its variants, such as PONOS-Net, PONOS-Weighted, and PONOS-Early.
6 Empirical Evaluation
This section provides empirical support for CSF: toxicity is not a fixed property of text in isolation, but a contextual social signal that is realised in interaction. Our definition decomposes toxicity into two components: (i) perceived norm violation (contradiction of accepted morality and norms of interaction) and (ii) the stress or disruption this perceived violation induces in an audience. Accordingly, we evaluate PONOS as a reception-based measure of the second component (stress or disruption manifested behaviourally), and we treat widely used text-only toxicity instruments as reference measures that primarily operationalise the first component (perceived norm violation) under assumed or averaged contexts.
We therefore design the evaluation as validity evidence for a reception-based axis, rather than as a leaderboard. Concretely, we test (i) whether the reception signal can be instantiated reliably in-domain, (ii) whether it is estimable without access to future replies, and (iii) whether it provides information that is distinct from intrinsic (text-only) instruments. We operationalise these three goals as the following evaluation questions:
  • Q1 (Operationalisation / instrument selection): Which sentiment model yields an in-domain sentiment function $s(\cdot)$ that best matches human judgements of reply sentiment in this community?
  • Q2 (Estimability at inference time): Can measured PONOS be predicted from the post text alone, and does explicitly modelling likely reactions improve estimation?
  • Q3 (Construct divergence / complementarity): How does reception-based measurement (PONOS) relate to intrinsic-toxicity instruments (text-only toxicity detectors)? Where do they agree, and where do they systematically diverge? Does PONOS provide a complementary axis aligned with CSF, or is it reducible to existing intrinsic scores?
6.1 Dataset Creation
We curate a dataset from the Reddit community r/BlackPeopleTwitter. We choose this community for two reasons: (i) its scale (5.8M members) and (ii) explicit encouragement of African-American Vernacular English (AAVE) and community-specific slang use. This setting is useful for studying context sensitivity, since dialectal and in-group expressions are frequently misinterpreted by generic toxicity systems trained on broad web data, effectively penalising speakers for communicating in their own dialect.
In total, we collected 401,931 posts containing 15,922,851 comments.
6.1.1 Human Annotation for Selecting $s(\cdot)$
To instantiate the sentiment function $s(\cdot)$ in Eq. 1, we construct a small, high-quality labelled set of posts and replies, since replies constitute the reception signal from which PONOS is computed.
We sample 56 user messages, each with at least four replies. If a message had too many replies, we retained only the top seven by score, resulting in a total of 273 replies.
Annotators and quality control.
We recruit 11 annotators via Prolific platform, all self-identifying as Black/African American U.S. nationals and regular internet users (7 men, 4 women; ages 21–60). We include attention checks and examine consistency. Four annotators show strong engagement and reliable performance; we retain their annotations for analysis.
Agreement and aggregation.
Inter-annotator agreement is Fleiss’ $\kappa=0.48$ with 67.8% pairwise agreement, consistent with moderate agreement for subjective judgement tasks. Final labels are obtained by majority vote; in rare 2-2 ties, an additional label is provided by a researcher following the same guidelines.
6.2 Experimental Setup
All experiments were run on NVIDIA H100 GPUs, totalling 883 compute hours. Model training and evaluation were performed using the February 2025 release of the unsloth library (unslothai2024). Metrics were computed using scikit-learn v1.5.0 scikit-learn.
Context provided to models.
For reply sentiment classification, we provide as much conversational context as feasible: the original comment, the reply text, the title of the post, and (when applicable) the set of replies under classification. We found that providing this context improved model’s performance. Prompts used with LLMs are provided in Appendix E.
6.3 Step 1 (Q1): In-Domain Instrument Selection for $s(\cdot)$
We evaluate multiple candidate sentiment models on the labelled reply set in order to select $s(\cdot)$ for Eq. 1. This step is best viewed as instrument selection: we choose the model that most reliably reproduces in-domain human judgements of negative reception, rather than optimising for generic sentiment benchmarks.
Table 1 reports performance for detecting the negative sentiment class on our labelled data.
Model Prec. Rec. F1
RoBERTa (samlowe2023roberta) 0.63 0.22 0.32
ModernBERT (cirimus2024modernbert) 0.71 0.21 0.32
Twitter-RoBERTa (barbieri-etal-2020-tweeteval) 0.58 0.64 0.61
GPT-4o (openai2024systemcard) 0.67 0.82 0.74
LLaMA 3 70B 0.81 0.70 0.75
Qwen 3 70B 0.76 0.80 0.78
Table 1: Precision, recall, and F1 scores for the negative sentiment class across candidate models on our labelled reply set.
Qwen 3 yang2025qwen3 achieved the strongest overall performance and is used as $s(\cdot)$ in all subsequent experiments. During error analysis, we found that most misclassifications occur in examples with low inter-annotator agreement. These instances often involved ambiguous language that led both assessors and models to varied interpretations.
6.4 Step 2 (Q2): Estimating PONOS from Post Text
With $s(\cdot)$ fixed, we compute PONOS from observed replies and then ask whether it can be accurately predicted at inference time (i.e. before replies are available).
Approach A: Direct PONOS Estimation
We train regression models to predict the measured PONOS value from the post text alone. This estimates reception without explicit reaction modelling and serves as a lightweight baseline suitable for low-latency settings.
Approach B: PONOS via Modelled Reactions
We also evaluate a reaction-modelling pipeline: a language model generates plausible replies to a post, and we apply $s(\cdot)$ to these generated replies to estimate PONOS. This directly instantiates CSF’s hypothesis that toxicity is reflected in likely community responses.
To separate the contributions of domain adaptation and reaction modelling, we evaluate two variants using LLaMA 3 70B llama2024herd: (i) reply prediction (RP) with PONOS inference via $s(\cdot)$, and (ii) domain-tuned reply prediction (DTRP).
6.4.1 Domain-Specific Pretraining Details
For DTRP, we pretrained LLaMA 3 70B on r/BlackPeopleTwitter data to better capture community style and norms. We used the unsloth framework to load and fine-tune the base model in 4-bit precision. The model was adapted using LoRA with rank 64 for efficient parameter updates. Gradient checkpointing and rslora were employed to further reduce memory overhead. All training and inference were conducted using bfloat16 precision and FlashAttention 2 to maximize memory efficiency.
Each training example consists of a Reddit comment and a prompt requesting five likely replies in the tone and style of r/BlackPeopleTwitter, separated by <|reply|>. The prompt template appears in Appendix E.
6.4.2 Step 3 (Q3): Comparing PONOS to Text-only Toxicity Scorers
We characterise how PONOS reception-based signal relates to intrinsic-toxicity instruments. To study divergence, we include intrinsic toxicity systems as reference instruments: the OpenAI Moderation API (omni-moderation-latest) (openai_moderation) and the Google Perspective API (jigsaw_perspective).
These systems estimate toxicity as a property of text and are therefore aligned with the perceived norm-violation component of our definition. In contrast, PONOS measures the reception or stress-disruption component via replies. We therefore focus on the relationship between these signals and on systematic regions of disagreement.
6.5 Results: Estimability and Construct Divergence
Estimating continuous PONOS.
Table 2 reports mean squared error (MSE) and mean absolute error (MAE) for continuous PONOS estimation. Two patterns emerge. First, strong text-only estimators (e.g. fine-tuned SentenceBERT) achieve competitive error, suggesting that some correlates of reception are recoverable from post text alone. Second, reaction modelling benefits substantially from domain adaptation: reply prediction without adaptation (RP) yields high error, while the domain-tuned reply prediction (DTRP) model attains the lowest error (MSE 0.042; MAE 0.152).
Together, these results support the claim that approximating community response can improve reception prediction when the generator is calibrated to community norms.
Model MSE MAE
Direct Prediction
TF–IDF + Ridge 0.046 0.173
SentenceBERT FT 0.043 0.166
Reply Prediction
LLaMA 3 8B RP 0.278 0.402
LLaMA 3 70B RP 0.138 0.304
GPT-4o RP 0.214 0.308
LLaMA 3 70B DTRP 0.042 0.152
Table 2: Mean squared error (MSE) and mean absolute error (MAE) for PONOS prediction. FT = fine-tuned model.
Construct relationship analysis: PONOS vs. intrinsic instruments.
We quantify the association between reception-based measurement and intrinsic-toxicity scoring using large-scale comparisons (n=88,717n=88{,}717) between measured PONOS and two widely deployed toxicity detection instruments: the OpenAI Moderation API and the Google Perspective API. For OpenAI, we summarise each post by the maximum category score ("max moderation score") and additionally report correlations for individual categories. For the Perspective API, we use the continuous toxicity score. [^1] In all cases, we report Spearman’s $\rho$ (monotonic association) and Pearson’s $r$ (linear association), with 95% confidence intervals via Fisher’s $z$ approximation.
Both instruments show positive but modest association with reception. For OpenAI, PONOS correlates with the max moderation score at $\rho=0.203$ (95% CI [0.197, 0.210]) and $r=0.195$ (95% CI [0.189, 0.201]); among categories, harassment shows the strongest monotonic association ($\rho=0.207$). For the Perspective toxicity score, the association is $\rho=0.176$ (95% CI [0.170, 0.183]) and $r=0.173$ (95% CI [0.167, 0.179]). Together, these results support the construct claim: intrinsic-toxicity scoring and reception-based PONOS are related but non-equivalent signals.
Finally, to test whether these patterns depend on a particular instrument, we measure inter-tool agreement. OpenAI max moderation score and Perspective toxicity correlate strongly ($\rho=0.866$, 95% CI [0.864, 0.868]), indicating that intrinsic instruments form a coherent axis that is weakly aligned with the reception-based PONOS axis.
Figure 1: Two-axis view of intrinsic toxicity (OpenAI max moderation score; x-axis) versus reception-based toxicity (PONOS; y-axis). Each quadrant reports prevalence (count and percentage; n=88,717n=88{,}717) and includes illustrative masked snippets.
Quadrant breakdown.
We discretise both axes into high versus low using quantile thresholds and count posts in each quadrant. Using top-25% thresholds on PONOS and max moderation score (Figure 1), we find that off-diagonal disagreement is substantial. For the OpenAI max moderation score, 14,513 posts (16.4%) are high-PONOS/low-moderation and 14,158 (16.0%) are low-PONOS/high-moderation (32.3% off-diagonal total). The Perspective API shows the same pattern: 15,174 posts (17.1%) are high-PONOS/low-intrinsic and 14,887 (16.8%) are low-PONOS/high-intrinsic (33.9% off-diagonal).
This supports the view that reception and intrinsic toxicity are related but non-equivalent axes.
To interpret the disagreement regions, we analyse two extreme divergence sets constructed from posts where reception and intrinsic instruments disagree most strongly. We emphasise that these are not universal "model failures"; rather, they are diagnostic cases that make construct mismatch concrete.
Reception-high / instrument-low: missed norm violations.
These posts elicit consistently hostile, aggressive, or sarcastic replies (maximal negative reception) yet receive low intrinsic scores. On average, the OpenAI max moderation score is 0.090 and the Perspective API toxicity score is 0.155.
These cases often involve sarcasm, pragmatic antagonism, or contextually coded norm violations that are legible to the community but lack strong lexical markers.
Reception-low / instrument-high: overflagging benign dialect and in-group language.
Conversely, these posts receive no negative reception but are flagged as highly toxic by intrinsic scorers. The OpenAI max moderation score averages 0.757 and the Perspective API toxicity score averages 0.674.
The content is dominated by dialectal expressions, humour, and identity-referential language that is benign in-community but resembles out-group toxic markers in generic training data.
Takeaway.
Across both extremes, divergence is systematic: lexically "clean" antagonism can provoke strong negative reception while remaining low-scoring intrinsically, whereas culturally situated profanity and in-group language can be intrinsically flagged while eliciting no negative reception. This provides concrete validity evidence for CSF’s two-axis view: reception-based measurement is complementary to intrinsic toxicity, and helps avoid both missed harms and false alarms in dialect- and context-rich communities.
7 Conclusion
We reframed toxicity as a contextual, interactional phenomenon that arises when a communicative act is perceived to violate local norms and thereby induces stress or disruption in its audience.
We proposed the Contextual Stress Framework (CSF) and operationalised a reception-based signal, PONOS, as the proportion of negative reactions to an utterance. Empirically, we provided validity evidence for this reception axis: we selected an in-domain sentiment instrument for $s(\cdot)$, we showed that PONOS is estimable at inference time (with domain-tuned reaction modelling performing best), and demonstrated systematic construct divergence from intrinsic toxicity instruments. In particular, large and interpretable off-diagonal regions show that reception-based measurement is complementary to intrinsic toxicity, and can help avoid both missed harms and false alarms in dialect- and context-rich communities.
Taken together, these findings support CSF’s central claim: intrinsic-toxicity scores and reception-based measurement are related but non-equivalent. They form complementary axes for analysing harm in context.
Future work can extend this approach by developing richer reception signals beyond reply polarity, validating CSF across platforms and cultures, incorporating multimodal context, and studying interventions that use both axes to mitigate harm.
8 Limitations
We acknowledge several important limitations of our work:
  • PONOS is not a universal or intrinsic measure of toxicity. It is a context-dependent, reception-driven proxy for perceived social stress. Rather than determining what a message is, PONOS captures how it is received by a particular audience in a particular setting.
  • As such, PONOS is best understood as a complementary signal that augments text-intrinsic toxicity models. It is most informative in settings where audience response is observable and sufficiently large. In Appendix C, we describe extensions such as PONOS-Weighted, PONOS-Net, and PONOS-Early, which account for user influence, network structure, and early-stage reception.
  • Stress is not synonymous with toxicity. Not all norm violations are harmful (e.g. satire, protest), and not all harm produces measurable stress. Our metric captures perceived disruption, not ethical correctness. It defines toxicity at the intersection of norm violation and stress response—not stress alone—to avoid conflating necessary discomfort with communicative harm; however, this separation may not always be obvious.
  • Sentiment classification is an imperfect proxy. PONOS relies on sentiment classifiers that are known to carry cultural, demographic, and linguistic biases. Their subjectivity introduces noise, especially in edge cases such as sarcasm or coded language.
  • Community-level perception $\neq$ moral truth. PONOS reflects what provokes disapproval in a particular group, not a universal judgement. What is toxic in one community may be normative in another. This is a strength—but also a limitation.
  • Susceptibility to manipulation. Like all crowd-based signals, PONOS can be gamed through brigading, reply injection, or sockpuppetry. Mitigating these risks may require extensions such as reply weighting (e.g. verified users, community trust scores, vulnerability-aware weighting) or time-window filters.
  • Blind spots. PONOS cannot capture silent harm—for example, when content is ignored, downvoted, or drives users away without leaving replies. It may miss damage caused by exclusion or indirect effects.
  • Not a moral compass. PONOS is not a substitute for ethical judgement. It models perception, not morality, and should be used alongside normative frameworks rather than in place of them.
9 Ethical Considerations
Any metric used in content moderation, filtering, or prioritisation carries normative weight. While we position CSF and PONOS as descriptive tools—reflecting community reactions rather than imposing moral standards—we recognise their potential for misuse or overreach.
Moderation decisions based solely on PONOS scores risk reinforcing existing biases or silencing dissent. High PONOS values should not automatically lead to suppression; rather, they should signal the need for human oversight, contextual analysis, and, where appropriate, dialogic intervention.
We explicitly do not claim that CSF and PONOS reflect objective truth. They reflect perceived social friction—a volatile signal that can be shaped by group norms, visibility effects, and adversarial influence.
Instead, we call for responsible deployment:
  • Use CSF and PONOS transparently and explainably.
  • Pair them with community input, human review, and ethical guidance.
  • Monitor for drift, manipulation, and misalignment with social values.
  • Consider their use not only for suppression, but also for norm diagnosis, echo-chamber detection, and tracking semantic change.
Ultimately, we advocate for a shift in perspective: from minimising toxicity to mapping social tension. This enables more pluralistic, adaptive, and humane approaches to language technology—aligned not with abstract ideals, but with the lived experience of language in context.
9.1 Data Collection Ethical Considerations
We complied with Reddit’s terms of service and did not collect sensitive personal information. Participants were informed about the task and that the data would be used for research. Annotators were compensated at \$12/hour in accordance with Prolific’s recommended payment guidelines. Full instructions are provided in Appendix D.
Appendix A Detailed Basis for CSF
To move beyond the limitations of current toxicity detection approaches, we synthesise insights from multiple disciplines to develop a framework for understanding verbal harm.
A.1 Personal level
The experience of verbal harm is fundamentally affective, rooted in how individuals perceive threats to personal and social stability. Research in psychology and neuroscience consistently shows that such harm does not arise from the intrinsic properties of language alone, but from context-dependent appraisals, emotional expectations, and violations of predictive frameworks.
In psychology, appraisal theory lazarus1991emotion posits that emotional and stress responses arise when individuals evaluate events as threatening to their goals, values, or social standing. From this perspective, communication becomes harmful not solely because of its content, but because it is perceived as violating significant personal concerns.
Research on social pain further demonstrates that symbolic threats can activate biological stress systems in much the same way as physical injury eisenberger2003does; lieberman2009neuroscience.
Neuroscientific findings indicate that the brain processes language in an anticipatory, predictive, and context-sensitive manner. Predictive coding models friston2010free; kleckner2017evidence; barrett2015interoceptive suggest that the brain continuously generates expectations about emotional and social states. When communication violates these expectations, it triggers a stress response.
Physiological indicators of stress, such as heart rate variability, are influenced by exposure to social threat shaffer2014healthy. Event-related potential studies ding2016emotion; liang2018contextual show that negative emotional language disrupts neural integration, while chronic exposure to social stressors is associated with pro-inflammatory genetic activity slavich2013emerging.
Together, this body of research demonstrates that responses to harmful or toxic speech are not merely psychological but also physiological.
A.2 Group level
Beyond direct victimisation, witnessing harm inflicted on others can also elicit stress responses, signalling potential threats within the social environment struiksma2022insulting.
Sociology and social psychology demonstrate that responses to verbal behaviour are governed by group norms, emotional expectations, and mechanisms of norm enforcement elster2007explaining. Research by Bicchieri_2005 and CIALDINI1991201 shows that social norms are maintained through collective emotional and behavioural reactions to deviance. Importantly, norm violations are not evaluated in isolation, but interpreted through the emotional dynamics of the group, which shape perceptions of harm.
The concepts of emotional collectives vankleef2015emotions and emotional norm-shaping processes VANKLEEF2024; sari2024role further emphasise how shared emotional responses reinforce or reshape social expectations in real time.
Moral Foundations Theory similarly highlights that emotional reactions to norm violations differ across groups and cultures, leading to divergent judgements about what constitutes harmful or toxic speech GRAHAM201355.
From this perspective, toxicity functions as a social signal: a disruption in the moral and emotional fabric of the group.
A.3 Cultural and Linguistic Level
Linguistic and sociolinguistic research demonstrates that meaning is co-constructed in interaction and shaped by contextual framing. Harmful speech cannot be reduced to individual words or expressions divorced from their social use.
Pragmatics clark1996using; Sperber1995-SPER emphasises that meaning is not fully encoded in words, but inferred through shared background knowledge and conversational context. An identical phrase may be benign in one setting and deeply offensive in another, depending on tone, power relations, and prior discourse.
Politeness and impoliteness theory Brown_Levinson_1987; CULPEPER1996349 similarly show that perceived rudeness, offence, or aggression are emergent properties of interaction rather than intrinsic features of utterances.
Language ideologies woolard1994language; irvine2001style further demonstrate that judgements about what counts as "toxic" or "neutral" language are shaped by power relations and dominant cultural narratives.
Historical linguistics and semantics highlight the instability of meaning across time and context allan2006forbidden; Traugott_Dasher_2001. Words such as "nice" (originally meaning "foolish") or "fag" (benign in British English but a slur in American usage) illustrate how emotional and moral associations shift across cultural and historical settings, undermining any fixed, universal standard of toxicity.
Finally, multilingualism and language variation introduce further complexity. Emotional resonance and perceived offensiveness vary across languages due to differing socialisation patterns Dewaele01062004. What is taboo or offensive in one language may carry no such connotation in another.
Taken together, these perspectives argue strongly against treating harmful language as a static or text-intrinsic phenomenon. Toxicity cannot be reliably detected by identifying “bad words” in isolation, but only by modelling how language interacts with contextual, social, and cultural expectations.
Appendix B Details on Understanding of Context
Context structures interpretation. It governs what is considered appropriate, offensive, or harmful within a given setting. To support operational clarity, we distinguish four key dimensions:
Verbal context refers to the surrounding linguistic material — what is said before, after, or alongside the utterance. It determines how meaning is disambiguated, particularly in cases of sarcasm, reappropriation, or coded language.
Situational context includes the physical, social, and institutional setting of communication: who is speaking, to whom, in what place, and under what conditions. Normative expectations can vary dramatically between settings — for example, language deemed acceptable at a music festival may be inappropriate in a courtroom or classroom. This dimension also encompasses digital and symbolic environments such as online platforms, workplace hierarchies, or subcultural spaces.
Cognitive context refers to the mental and emotional states of the communicative participants — including intentions, assumptions, prior knowledge, and emotional stance. Many speech acts, such as jokes, flirtation, or irony, can only be correctly interpreted when cognitive context is inferred or shared.
Cultural context involves the broader moral frameworks, communicative conventions, and ideological norms that vary across communities. A behaviour considered respectful in one culture may be viewed as inappropriate or even threatening in another. For example, direct individualistic expression is normative in many Western contexts, but may be perceived as disruptive or disrespectful in collectivist settings.
Given this multidimensional variability, any model of toxicity detection must be contextually grounded and culturally adaptive. Static or universalist rule sets will inevitably misclassify utterances as they fail to consider how meaning emerges within interaction.
Appendix C PONOS Variants
C.1 Extensions and Variants
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context.
PONOS-Net.
We define the net sentiment variant as:
$\text{PONOS}_{\text{net}}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}s(r_{i})$ (2)
This version captures the overall affective balance of replies, distinguishing polarised threads (with both positive and negative responses) from uniformly negative ones.
PONOS-Weighted.
To account for user-level features, such as author credibility, social influence, or vulnerability - we define a weighted version:
$\begin{split}\text{PONOS}_{\text{w}}(x)=\frac{1}{\sum_{i}w(r_{i})}\sum_{r_{i}\in R(x)}w(r_{i})\cdot\mathbb{I}[s(r_{i})=-1]\end{split}$ (3)
Here, $w(r_{i})$ is a weight assigned to each reply, reflecting community-specific salience (e.g., upvotes, karma, user type).
PONOS-Early.
To mitigate feedback loop effects (e.g., reply priming), we define an early-window variant that considers only the first $k\ll n$ replies:
$\text{PONOS}_{\text{early}}(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}[s(r_{i})=-1]$ (4)
This variant reflects the immediate, unprimed perception of a post, prior to potential community escalation or reinforcement.
C.2 Implementation Considerations
The sentiment function $s(\cdot)$ can be realised through multiple approaches:
  • Manual annotation by culturally aligned raters;
  • Off-the-shelf sentiment classifiers (e.g., RoBERTa, BERTweet, GPT-based);
  • Community-specific fine-tuned models.
We emphasize that the interpretability and validity of PONOS depend on the alignment between the sentiment model and the communicative norms of the target community. In our experiments, we rely on human-annotated sentiment labels sourced from subreddit-native annotators to ensure cultural fidelity.
Appendix D Assessors’ Instructions
Dear Participant,
Thank you for taking the time to participate in our survey. Your input is invaluable to us. In this survey, you will be asked to classify reactions expressed in replies to comments from discussions on the Reddit community.
Task:
- Given a comment and its replies from a discussion on the Reddit community r/BlackPeopleTwitter, classify the reaction expressed in each reply as one of the following:
- "approval (comment)" - Agreement/support of the comment itself.
- "approval (subject)" - Agreement/support of the subject mentioned in the comment.
- "neutral" - No strong reaction for or against the comment or subject.
- "condemnation (comment)" - Disapproval of the comment itself.
- "condemnation (subject)" - Disapproval of the subject mentioned in the comment.
- Be sure to identify whether approval/condemnation is directed at the comment or the subject.
Examples:
Topic: "New office policy - no phones allowed"
Comment: "The new policy is a disaster, its making everything worse."
1. Reply: "I totally agree, its a complete mess."
- Classification: approval (comment)
2. Reply: "I think we should give it some time to see the effects."
- Classification: neutral
3. Reply: "This policy is the worst decision ever made by this administration."
- Classification: condemnation (subject)
4. Reply: "I dont have a strong opinion on this."
- Classification: neutral
5. Reply: "LOL, shut up Karen, nobody cares what you think."
- Classification: condemnation (comment)
6. Reply: "I kinda like the new policy."
- Classification: approval (subject)
By following these instructions, you will help us understand the community dynamics and how different reactions are expressed within the r/BlackPeopleTwitter community.
Your participation is crucial for the success of this research. Thank you!
Appendix E Prompts
Sentiment classification prompt:
You are a sentiment analysis assistant trained to understand African American Vernacular English (AAVE) and online Black community discourse.
Your task is to classify the emotional tone of REPLY message as Positive, Neutral, or Negative.
Your analysis supports a larger framework for detecting *perceived toxicity* - understanding how it might cause stress or social tension within the community.
Use all available context:
- Topic: the post that started the thread
- Comment: the message being replied to
- Reply: the specific message to analyze
Consider slang, tone, cultural style, and conversational intent.
- Positive: Agreement, Pride, joy, affirmation, playfulness, flirting, or supportive sarcasm. Being angry TOGETHER against something also goes here.
- Neutral: Factual comments, observations, dry or ambiguous tone, rhetorical questions.
- Negative: Disagreement, confrontation, arguing, complaints, passive aggression, warnings, critique, or social "shade."
Always output one of: Positive’, Neutral’, or Negative without explanation. Be sure to not miss any Negative’.
Format your answer as "ANSWER:"
FEW-SHOT EXAMPLES:
Topic: Trump won
Comment:" I think this is ok as long as we are not getting democrat president
Reply: Nigga this is a HILARIOUS take
ANSWER: Negative
Topic: New record lol
Comment: That clip had me dying, fr.
Reply: Ayo I been laughing for 10 minutes, why nobody brings up that he claimed to be THE MASTER
ANSWER: Positive
Topic: Gotta do whacha gotta do
Comment: Once I was texting with girl for 6 months to get a picture from her
Reply: Bro, Ive ended up writing her real letters. desparate shit
ANSWER: Neutral
Topic: Economy is fucked
Comment: I finna strat looting just to get food on the table
Reply: they are making us do it
ANSWER: Positive
Reaction prediction prompt:
{"role": "system", "content": "You are a culturally aware assistant who creates authentic replies from r/BlackPeopleTwitter."},
{"role": "user", "content": (
"Below is a Reddit comment. Generate the top 5 likely replies from r/BlackPeopleTwitter, "
[…]